# Acceleration as a circular motion along an imaginary circle: Kubo-Martin-Schwinger condition for accelerating field theories in imaginary-time formalism

### v2

@@ 8 KMS relations in Rindler coordinates
$\theta_{(j)}=\theta+j\alpha,\qquad\zeta_{(j)}=\zeta,\qquad j\in{\mathbb{Z}}\,.$ (39)
@@ 9 Energy-momentum tensor with the accelerated KMS conditions
$\Theta_{\mu\nu}=-\frac{1}{2}\lim_{x^{\prime}\rightarrow x}{\rm tr}[\gamma^{E}_{\mu}(\partial_{\nu}-\partial_{\nu^{\prime}})\Delta S^{(\alpha)}_{E}].$ (43)
@@
$\Theta_{\mu\nu}=\sum_{j\neq 0}\frac{(-1)^{j}}{\pi^{2}}\left[\delta_{\mu\lambda}\cos\tfrac{j\alpha}{2}+\left(\delta_{\mu 0}\delta_{\lambda 3}-\delta_{\mu 3}\delta_{\lambda 0}\right)\sin\tfrac{j\alpha}{2}\right]\\ \times\left[\frac{R_{\nu\lambda}^{(j)}+\delta_{\nu\lambda}}{\Delta X^{4}_{(j)}}-\frac{4\Delta X^{(j)}_{\lambda}}{\Delta X_{(j)}^{6}}(R^{(j)}_{\nu\kappa}+\delta_{\nu\kappa})\Delta X^{(j)}_{\kappa}\right].$ (48)
@@
$\mathcal{E}_{\xi}$ $=\frac{3[\alpha T(x)]^{4}}{16\pi^{2}}\left[G_{4}(\alpha)+4\xi G_{2}(\alpha)\right],$
$\mathcal{P}_{\xi}$ $=\frac{[\alpha T(x)^{4}]}{16\pi^{2}}\left[G_{4}(\alpha)+\frac{4}{3}\left(1-3\xi\right)G_{2}(\alpha)\right],$
$\pi_{s}^{\xi}$ $=-\frac{[\alpha T(x)]^{4}}{12\pi^{2}}(1-6\xi)G_{2}(\alpha),$ (51)
@@
Substituting now $G_{n}(\alpha)={\rm Re}[i^{-n}\widetilde{G}_{n}(i\beta_{T}\phi)\rfloor_{\phi\rightarrow-ia}]$ into Eq. (51) gives Eq. (10) for the conformal coupling $\xi=1/6$. For minimal coupling $\xi=0$ or a generic non-conformal coupling $\xi\neq 1/6$, we recover the results of Refs. [37, 54].
@@ 10 Fractalization of thermodynamics
$\langle\phi^{2}\rangle^{(p,q)}$ $=\frac{[\alpha T(x)]^{2}}{96\pi^{2}}(q^{2}-1),$ (57a)
$\mathcal{E}_{\xi}^{(p,q)}$ $=\frac{[\alpha T(x)]^{4}}{480\pi^{2}}(q^{2}-1)(q^{2}+11+60\xi),$ (57b)
$\mathcal{P}_{\xi}^{(p,q)}$ $=\frac{[\alpha T(x)]^{4}}{1440\pi^{2}}(q^{2}-1)(q^{2}+31-60\xi),$ (57c)
$\pi_{s;\xi}^{(p,q)}$ $=-\frac{[\alpha T(x)]^{4}}{72\pi^{2}}(1-6\xi)(q^{2}-1),$ (57d)
@@ Acknowledgements
This work is supported by the European Union - NextGenerationEU through the grant No. 760079/23.05.2023, funded by the Romanian ministry of research, innovation and digitalization through Romania’s National Recovery and Resilience Plan, call no. PNRR-III-C9-2022-I8.

### v3

@@ 8 KMS relations in Rindler coordinates
$\theta_{(j)}=\theta-j\alpha,\qquad\zeta_{(j)}=\zeta,\qquad j\in{\mathbb{Z}}\,.$ (39)
@@ 9 Energy-momentum tensor with the accelerated KMS conditions
$\Theta_{\mu\nu}=\frac{1}{2}\lim_{x^{\prime}\rightarrow x}{\rm tr}[\gamma^{E}_{\mu}(\partial_{\nu}-\partial_{\nu^{\prime}})\Delta S^{(\alpha)}_{E}].$ (43)
@@
$\Theta_{\mu\nu}=-\sum_{j\neq 0}\frac{(-1)^{j}}{\pi^{2}}\left[\delta_{\mu\lambda}\cos\tfrac{j\alpha}{2}+\left(\delta_{\mu 0}\delta_{\lambda 3}-\delta_{\mu 3}\delta_{\lambda 0}\right)\sin\tfrac{j\alpha}{2}\right]\\ \times\left[\frac{R_{\nu\lambda}^{(j)}+\delta_{\nu\lambda}}{\Delta X^{4}_{(j)}}-\frac{4\Delta X^{(j)}_{\lambda}}{\Delta X_{(j)}^{6}}(R^{(j)}_{\nu\kappa}+\delta_{\nu\kappa})\Delta X^{(j)}_{\kappa}\right].$ (48)
@@
$\mathcal{E}_{\xi}$ $=\frac{3[\alpha T(x)]^{4}}{16\pi^{2}}\left[G_{4}(\alpha)-4\xi G_{2}(\alpha)\right],$
$\mathcal{P}_{\xi}$ $=\frac{[\alpha T(x)^{4}]}{16\pi^{2}}\left[G_{4}(\alpha)-\frac{4}{3}\left(1-3\xi\right)G_{2}(\alpha)\right],$
$\pi_{s}^{\xi}$ $=\frac{[\alpha T(x)]^{4}}{12\pi^{2}}(1-6\xi)G_{2}(\alpha),$ (51)
@@
Substituting now $G_{n}(\alpha)={\rm Re}[i^{-n}\widetilde{G}_{n}(\beta_{T}\phi)\rfloor_{\phi\rightarrow-ia}]$ into Eq. (51) gives Eq. (10) for the conformal coupling $\xi=1/6$. For minimal coupling $\xi=0$ or a generic non-conformal coupling $\xi\neq 1/6$, we recover the results of Refs. [37, 54].
@@ 10 Fractalization of thermodynamics
$\langle\phi^{2}\rangle^{(p,q)}$ $=\frac{[\alpha T(x)]^{2}}{48\pi^{2}}(q^{2}-1),$ (57a)
$\mathcal{E}_{\xi}^{(p,q)}$ $=\frac{[\alpha T(x)]^{4}}{480\pi^{2}}(q^{2}-1)(q^{2}+11-60\xi),$ (57b)
$\mathcal{P}_{\xi}^{(p,q)}$ $=\frac{[\alpha T(x)]^{4}}{1440\pi^{2}}(q^{2}-1)(q^{2}-9+60\xi),$ (57c)
$\pi_{s;\xi}^{(p,q)}$ $=\frac{[\alpha T(x)]^{4}}{72\pi^{2}}(1-6\xi)(q^{2}-1),$ (57d)
@@ Acknowledgements
The authors are indebted to our colleagues: Dr. Andrea Palermo, Dr. Matteo Buzzegoli and Prof. Francesco Becattini for carefully reading our work and for drawing our attention to (some of) the errors in previous versions of this paper. This work is supported by the European Union - NextGenerationEU through the grant No. 760079/23.05.2023, funded by the Romanian ministry of research, innovation and digitalization through Romania’s National Recovery and Resilience Plan, call no. PNRR-III-C9-2022-I8.